# Beyond Sharp Minima: Robust LLM Unlearning via Feedback-Guided Multi-Point Optimization

### v1

@@ 5.2 Experiment Results
Table 1: Evaluation of Unlearning Effectiveness and Model Utility of on MUSE News and MUSE Books, evaluated under two unlearning settings: LLaMA2-7B on News and ICLM-7B on Books.
@@
In Figure 5(a), we demonstrate the unlearning performance of five distinct $\mathcal{MU}$ methods integrated with the StableUN on WMDP, evaluated against input-level adversarial prompts generated by enhanced GCG (Łucki et al., 2025). It can be observed that StableUN exhibits a significant effect in suppressing the recovery of unlearning performance caused by jailbreak attacks; specifically, our method achieves an average improvement of 14.4%. This is attributed to the smoother loss landscape induced by the adversarial and randomized perturbations employed in this work, as such smoothing effects are known to aid in defending against input-level adversarial attacks (Shang & Wei, 2025; Robey et al., 2023). We also provide generated examples of NPO and GA under jailbreak attacks in Appendix A.7. In Figure 5(b) and 5(c), we plot the KL divergence of each output token between the unlearned model $f_{U}$ (GA/NPO) and the original model $f_{O}$. A higher KL divergence indicates more effective unlearning. Our method demonstrates a larger KL divergence, which mitigates the shallow optimization problem (Pu et al., 2025) in unlearning and enhances the robustness against jailbreak attacks.
@@
Figure 5: Evaluations under jailbreak attacks: (a) $\mathcal{MU}$ effectiveness comparison with feedback guidance added. (b)(c) KL divergence for output token between $f_{U}$ and the $f_{O}$ for GA and NPO.

### v3

@@ 5.2 Experiment Results
Table 1: Evaluation of Unlearning Effectiveness and Model Utility on MUSE News and MUSE Books, evaluated under two unlearning settings: LLaMA2-7B on News and ICLM-7B on Books.
@@
In Figure 5(a), we demonstrate the unlearning performance of five distinct $\mathcal{MU}$ methods integrated with the StableUN on WMDP, evaluated against input-level adversarial prompts generated by enhanced GCG (Łucki et al., 2025). It can be observed that StableUN exhibits a significant effect in suppressing the recovery of unlearning performance caused by jailbreak attacks; specifically, our method achieves an average improvement of 14.4%. This is attributed to the smoother loss landscape induced by the adversarial and randomized perturbations employed in this work, as such smoothing effects are known to aid in defending against input-level adversarial attacks (Shang & Wei, 2025; Robey et al., 2023). We also provide generated examples of NPO and GA under jailbreak attacks in Appendix A.7. In Figure 5(b) and 5(c), we plot the KL divergence of each output token between the unlearned model $f_{\text{U}}$ (GA/NPO) and the original model $f_{\text{O}}$. A higher KL divergence indicates more effective unlearning. Our method demonstrates a larger KL divergence, which mitigates the shallow optimization problem (Pu et al., 2025) in unlearning and enhances the robustness against jailbreak attacks.
@@
Figure 5: Evaluations under jailbreak attacks: (a) $\mathcal{MU}$ effectiveness comparison with StableUN added. (b)(c) KL divergence for output token between $f_{\text{U}}$ and $f_{\text{O}}$ for GA and NPO, respectively.